OpenStudy (anonymous):

–2(y – 3) = 2y – 6 whats y?

OpenStudy (anonymous):

distribute -2 to get... -2y+6=2y-6 then put like terms together 6=4y-6 12=4y divide 4 from both sides to get your answer 12/4=y y=3